### v1

# Simulation of the magnetic Ginzburg-Landau equation via vortex tracking

Abstract.
This paper deals with the numerical simulation of the 2D magnetic time-dependent Ginzburg-Landau (TDGL) equations in the regime of small but finite (inverse) Ginzburg-Landau parameter $\epsilon$ and constant (order $1$ in $\epsilon$) applied magnetic field. In this regime, a well-known feature of the TDGL equation is the appearance of quantized vortices with core size of order $\epsilon$. Moreover, in the singular limit $\epsilon\searrow 0$, these vortices evolve according to an explicit ODE system. In this work, we first introduce a new numerical method for the numerical integration of this limiting ODE system, which requires to solve a linear second order PDE at each time step. We also provide a rigorous theoretical justification for this method that applies to a general class of 2D domains. We then develop and analyze a numerical strategy based on the finite-dimensional ODE system to efficiently simulate the infinite-dimensional TDGL equations in the presence of a constant external magnetic field and for small, but finite, $\epsilon$. This method allows us to avoid resolving the $\epsilon$-scale when solving the TDGL equations, where small values of $\epsilon$ typically require very fine meshes and time steps. We provide numerical examples on a few test cases and justify the accuracy of the method with numerical investigations.
Key words and phrases:
Ginzburg-Landau, magnetic vortices, Schrödinger flow, superconductor model
2020 Mathematics Subject Classification:
Primary: 65M12, 35A21 Secondary 35Q55
Funding information: DFG – Project-ID 442047500 – SFB 1481.
©2025 by the authors. Faithful reproduction of this article, in its entirety, by any means is permitted for noncommercial purposes.
1. Introduction
1.1. Motivation
The Ginzburg-Landau (GL) theory has played a central role in the description of superconductivity since its introduction by Ginzburg and Landau [GL50]. By encoding macroscopic quantum behavior through a complex order parameter, the GL framework bridges microscopic physics to observable phenomena such as permanent superconducting currents, the expulsion of external magnetic fields from the superconducting sample (Meissner effect), and the different phase transitions between superconducting, mixed, and normal states. Its time-dependent extension, the time-dependent Ginzburg–Landau (TDGL) equation, was later developed to describe nonequilibrium processes in superconductors [GE68, Sch66]. Moreover, while originally proposed on phenomenological grounds, the GL theory was later justified from the microscopic Bardeen-Cooper-Schrieffer (BCS) theory [BCS57, Gor59, FHSS12]. However, we note that a rigorous derivation of BCS theory from many-body quantum mechanics remains a major challenge in mathematical physics.
A particularly striking feature of the GL framework is the emergence of quantized vortices when the intensity of the applied magnetic field reaches a critical threshold [Abr57]. In superconductors, these vortices govern dissipation through flux motion and pinning, thereby determining critical currents and resistance under applied fields [BS65, Tin96, ST11]. Therefore, their dynamics, interactions, nucleation, and annihilation are central to understanding the dynamical response of type-II superconductors in technological applications. Beyond its original setting, GL-type models have also found applications in the study of Bose–Einstein condensates, where they serve as effective mean-field descriptions of coherent matter waves (see, e.g., [Aft06] for a mathematical overview). As such, the accurate simulation of vortex dynamics within the TDGL equation is of both fundamental and applied importance.
From a computational perspective, however, the resolution of vortex structures poses significant challenges. The characteristic vortex core size is set by the (inverse) Ginzburg–Landau parameter $\epsilon$, which may be very small in physically relevant regimes, e.g., for superconductors of type II. Standard finite element discretizations, while powerful in handling complex geometries, require extremely fine meshes to capture the core structure when $\epsilon\ll 1$, leading to prohibitively large computational costs even in the static case [DGP92, DH24, BDH25, DDH25]. This difficulty is amplified in time-dependent simulations, where vortices may move and interact dynamically. Despite the central importance of vortex dynamics, relatively few numerical methods are designed specifically for the small-$\epsilon$ regime (see [CKMS25] and references therein) and the development of efficient and accurate approaches for TDGL simulations in this setting remains an open and difficult problem [Du97, Du05, Bar05a, Bar05b, BMO11]. In the specific regime of a constant external magnetic field $h_{\rm ex}=\mathcal{O}(1)$ and $\epsilon\searrow 0$, the vortices move according to a finite dimensional ODE system (see for instance [KS11] for a rigorous derivation of this limiting system, and references therein). Extending our previous work [CKMS25] in the simpler case without magnetic field, we present in this paper a novel numerical method to tackle the simulation of the TDGL equations in the regime of small, but finite, $\epsilon$ under a constant external magnetic field. This method is designed to achieve higher accuracy for smaller values of $\epsilon$.
1.2. Main contribution
The main contributions of this article can be summarized as follows:
  1. (i)
    We propose and implement a numerical method to compute approximate solutions of the time-dependent Ginzburg-Landau equation with magnetic fields in the regime of small but finite $\epsilon$. This method builds on rigorous analytical results concerning the dynamics of the PDE solutions in the limit as $\epsilon\searrow 0$. In particular, it avoids the use of extremely fine meshes that are required by standard FEM methods to resolve the vortex core.
  2. (ii)
    We propose and implement a numerical scheme to efficiently simulate the vortex dynamics in the limit $\epsilon\searrow 0$. Moreover, we provide a theoretical justification for this scheme that applies to general Lipschitz domains.
1.3. Outline of the paper
The paper is structured as follows. In the next section, we introduce the necessary notations for precisely stating the TDGL equations and highlight some of its basic properties. We then recall some rigorous results on the asymptotic behavior of solutions to the TDGL equations in the limit where the inverse Ginzburg-Landau parameter $\epsilon$ vanishes and the applied magnetic field is kept constant. In Section 3, we describe the numerical scheme for simulating the asymptotic dynamics of vortices. In addition, we provide a theoretical justification for this numerical scheme, present numerical simulations on the vortex trajectories, and numerically analyze the convergence rates with respect to the discretization parameters. In Section 4, we present the numerical scheme to efficiently approximate solutions of the TDGL equation in the regime of small but finite $\epsilon$. We then illustrate our method with some numerical experiments. Finally, we compare the results with reference solutions computed via finite elements to numerically justify the accuracy of our method.
2. The time-dependent magnetic Ginzburg-Landau equation
Let us now introduce the necessary notations to state the time-dependent Ginzburg-Landau (TDGL) equations and recall a few basic facts about these equations. We then recall some important theoretical results regarding the asymptotic behavior of the solutions in the regime of interest to us. These results form the basis for our numerical method, which will be described in Section 4.
Throughout this paper, we let $\Omega\subset\mathbb{R}^{2}$ be an open bounded and simply connected subset with smooth boundary. In the phenomenological description of superconductors via the Ginzburg-Landau theory, there are two main variables of interest: the so-called order parameter $u:\Omega\rightarrow\mathbb{C}$, describing the density of superconducting Cooper pairs, and the associated magnetic vector potential $A:\Omega\rightarrow\mathbb{R}^{2}$ inside the superconducting sample $\Omega$. The Ginzburg-Landau energy of the pair $(u,A)$ then reads
$GL_{\epsilon}(u,A)=\frac{1}{2}\int_{\Omega}|\nabla_{A}u(x)|^{2}+|\operatorname{curl}A(x)-h_{\rm ex}|^{2}+\frac{(1-|u(x)|^{2})^{2}}{2\epsilon^{2}}\mathrm{d}x,$ (2.1)
where $\nabla_{A}u\coloneqq\nabla u-\imath Au$ is the covariant gradient, $\operatorname{curl}A\coloneqq\partial_{1}A_{2}-\partial_{2}A_{1}$ is the rotational of $A$ (a real number since the domain is two dimensional), and $h_{\rm ex}\in\mathbb{R}$ is the applied external magnetic field. Typically, $h_{\rm ex}$ is assumed constant throughout the sample $\Omega$.
The time evolution of the order parameter and magnetic field can then be modeled (see [GE68, Sch66, Dor92]) via the following time-dependent Ginzburg-Landau equations:
$\delta_{\epsilon}\left(\partial_{t}u+\imath\Phi u\right)=-\nabla_{u}GL_{\epsilon}(u,A)=\nabla_{A}^{2}u+\frac{1}{\epsilon^{2}}u(1-|u|^{2}),$ (2.2)
$\sigma_{\epsilon}\left(\partial_{t}A+\nabla\Phi\right)=-\nabla_{A}GL_{\epsilon}(u,A)=-\operatorname{curl}h+(\imath u,\nabla_{A}u),$ (2.3)
where $h=\operatorname{curl}A$ is the magnetic field induced by $A$, $(u,v)=\mathrm{Re}(u\overline{v})=\frac{u\overline{v}+\overline{u}v}{2}$ is the standard inner-product on $\mathbb{R}^{2}\simeq\mathbb{C}$, $\operatorname{curl}h=-\nabla^{\perp}h=(\partial_{2}h,-\partial_{1}h)^{T}$ is the curl of the scalar magnetic field, $\delta_{\epsilon}=\alpha_{\epsilon}+\imath\beta_{\epsilon}$ is a complex relaxation parameter and $\sigma_{\epsilon}>0$ is the conductivity. In addition, we impose what is called the natural boundary conditions [^1] The term natural boundary conditions refers to the fact that Neumann boundary conditions are automatically satisfied by minimizers of the GL energy $GL_{\epsilon}(u,A)$ without any boundary constraints, i.e., in the spaces defined in (2.7)-(2.8).
$\nu\bm{\cdot}\nabla_{A}u=0,\quad\nu\bm{\cdot}(\partial_{t}A+\nabla\Phi)=0,\quad\mbox{and}\quad h=h_{\rm ex}\quad\mbox{on $\partial\Omega$.}$ (2.4)
The parameter $\epsilon>0$ is a material constant. In this work, we are interested in the regime of small $\epsilon>0$, which corresponds to superconductors of type II.
At first sight, it may seem that the electric potential $\Phi$ is an additional variable of the problem. However, note that by taking the divergence of (2.3), the electric potential is, up to a constant, uniquely determined as the solution of the following Poisson equation with Neumann boundary conditions:
$\Delta\Phi=\frac{1}{\sigma_{\epsilon}}\mathrm{div}\,j_{A}(u)\quad\mbox{in $\Omega$},\quad\partial_{\nu}\Phi=\frac{1}{\sigma_{\epsilon}}j_{A}(u)\bm{\cdot}\nu\quad\mbox{in $\partial\Omega$.}$ (2.5)
Moreover, note that the GL energy, and therefore the space of solutions to the GL equations, is invariant under the gauge transformation
$u\mapsto u\mathrm{e}^{\imath t\chi},\quad A\mapsto A+\nabla\chi,\quad\Phi\mapsto\Phi-\partial_{t}\chi,$ (2.6)
for a (sufficiently regular) function $\chi:\mathbb{R}_{+}\times\Omega\rightarrow\mathbb{R}$. Nevertheless, note that all of the physical quantities of interest for the model, namely the density $|u|^{2}$, the super-current $j_{A}(u)=(\imath u,\nabla_{A}u)$, the induced electric field $E=\partial_{t}A+\nabla\Phi$, and the induced magnetic field $h=\mathrm{curl}\,A$, are invariant under this gauge transformation. In particular, we are free to choose the most convenient gauge for calculations. Following [Spi03, KS11], we shall work with the Coulomb gauge here, i.e., we impose the additional conditions:
$\mathrm{div}\,A=0\quad\mbox{in $\Omega$,}\quad A\bm{\cdot}\nu=0\quad\mbox{in $\partial\Omega$}\quad\mbox{and}\quad\int_{\Omega}\Phi=0.$
Under these conditions, equation (2.2) is globally well-posed in the energy space, see e.g., [KS11, Remark 1.3] and [TW95].
2.1. Asymptotic regime and limiting dynamics
In this paper, we are mostly interested in the regime of small $\epsilon>0$. In this regime, one typically observes the formation of so-called vortices, which correspond to isolated zeros of the density $|u|^{2}$ with a non-zero winding number (or degree). A central topic in the study of superconductors is the description of the dynamics of such vortices. In particular, there are several works dedicated to deriving and justifying such dynamics with various degrees of mathematical rigor, see, e.g. [PR93, E94, TW95, LD97, OS98, GS06, SS04, SS07] and references therein. In this section, we recall the results from [Spi02, Spi03, KS11], which apply to the model equation studied here.
To this end, let us introduce the so-called renormalized energy which corresponds to the next-order contribution to the energy for minimizers of the GL energy with a given vortex configuration. Precisely, let $n\in\mathbb{N}$ and let $\Omega^{n}_{\ast}$ be the set
$\Omega_{\ast}^{n}\coloneqq\{a=(a_{1},a_{2},..,a_{n})\in\Omega^{n}:a_{j}\neq a_{i}\quad\mbox{for any $i\neq j$.}\}.$
Then, we can define the renormalized energy of a given vortex configuration $(a,d)\in\Omega_{\ast}^{n}\times\mathbb{Z}^{d}$ as
$W_{\Omega}(a,d;h_{\rm ex})=\lim_{\rho\searrow 0}\inf_{\begin{subarray}{c}u\in\mathcal{H}^{1}_{\rho}(a)\\ A\in\mathrm{H}^{1}_{\rm div}(\Omega)\end{subarray}}\left\{\frac{1}{2}\int_{\Omega_{\rho}(a)}|\nabla_{A}u|^{2}+|\mathrm{curl}\,A-h_{\rm ex}|^{2}\mathrm{d}x\right\}-\pi\sum_{j=1}^{n}d_{j}^{2}\log\frac{1}{\rho}$
where $\Omega_{\rho}(a)\coloneqq\Omega\setminus\cup_{j=1}^{n}B_{\rho}(a_{j})$,
$\mathrm{H}^{1}_{\mathrm{div}}(\Omega)\coloneqq\{A\in\mathrm{H}^{1}(\Omega;\mathbb{R}^{2}):\mathrm{div}A=0\mbox{ on $\Omega$,}\quad A\bm{\cdot}\nu=0\mbox{ on $\partial\Omega$}\},$ (2.7)
and
$\mathcal{H}^{1}_{\rho}(a)\coloneqq\{u\in\mathrm{H}^{1}(\Omega_{\rho}(a);\mathbb{S}^{1}):\quad\mathrm{deg}(u,a_{j})=d_{j}\quad\mbox{for any $j=1,...,n$}\}.$ (2.8)
Here, $\mathrm{deg}(u,a_{j})$ denotes the topological degree of $u$ around $a_{j}$,
$\mathrm{deg}(u,a_{j})\coloneq\frac{1}{2\pi}\int_{\partial B_{\rho}(a_{j})}\frac{1}{|u|^{2}}(\imath u,\partial_{\tau}u),$
which is well-defined for any non-vanishing $u\in\mathrm{H}^{\frac{1}{2}}(\partial B_{\rho}(a_{j});\mathbb{C})$ (see, e.g., [BM21]). By the results in [Spi03, Appendix A.1] (see also [LD97, Section 4.1]), the renormalized energy can be written more explicitly [^2] Here we adopt the sign convention for $\Xi$ in [KS11]. Note that this is the opposite sign convention from [Spi03, LD97]; in particular, the terms depending on $\Xi$ in the renormalized energy appear with the opposite sign there. as
$W_{\Omega}(a,d;h_{\rm ex})=-\pi\sum_{i\neq j}d_{i}d_{j}\log|a_{i}-a_{j}|-\pi\sum_{j=1}^{n}d_{j}\left(R(a_{j})+\Xi(a_{j})\right)+\frac{h_{\rm ex}^{2}}{2}|\Omega|+\frac{h_{\rm ex}}{2}\int_{\partial\Omega}\partial_{\nu}\Xi,$ (2.9)
where $R$ is the harmonic function satisfying
$\Delta R=0\quad\mbox{in $\Omega$}\quad\mbox{and}\quad R(x)=-\sum_{j=1}^{n}d_{j}\log|x-a_{j}|,\quad x\in\partial\Omega,$ (2.10)
and $\Xi$ solves
$\Delta^{2}\Xi-\Delta\Xi=2\pi\sum_{j=1}^{n}d_{j}\delta_{a_{j}}\quad\mbox{in $\Omega$,}\quad\Xi=0\quad\mbox{on $\partial\Omega$,}\quad\mbox{and}\quad\Delta\Xi=-h_{\rm ex}\quad\mbox{on $\partial\Omega$.}$ (2.11)
Moreover, let us introduce the constant $\gamma_{0}>0$ defined (see [BBH94]) as
$\gamma_{0}\coloneqq\lim_{r\rightarrow\infty}I(r,\epsilon)-\pi\log\frac{r}{\epsilon},$ (2.12)
where
$I(r,\epsilon)=\inf\left\{\frac{1}{2}\int_{B_{r}}|\nabla u|^{2}+\frac{(1-|u|^{2})^{2}}{2\epsilon^{2}}:u\in H^{1}(B_{r}),u=\mathrm{e}^{\imath\theta}\mbox{on $\partial B_{r}$}\right\},$
and define the gauged Jacobian as
$J_{A}(u)\coloneqq\frac{1}{2}\operatorname{curl}j_{A}(u)+\frac{1}{2}\operatorname{curl}A.$ (2.13)
We can now introduce the following definition of well-prepared states, that is, a sequence of $\epsilon$-dependent states whose energy and vorticity is compatible with a given vortex configuration.
Definition 2.1 (Well-prepared states).
Let $\{(u_{\epsilon},A_{\epsilon})\}_{\epsilon>0}\subset H^{1}(\Omega)\times H^{1}(\Omega;\mathbb{R}^{2})$ be a sequence of functions. Then we say that $\{(u_{\epsilon},A_{\epsilon})\}_{\epsilon>0}$ are well-prepared with respect to $a\in\Omega^{n}_{\ast}$ and $d\in\{\pm 1\}^{n}$ for some $n\in\mathbb{N}$ if it holds that
$GL_{\epsilon}(u_{\epsilon},A_{\epsilon};h_{\rm ex})=n\left(\pi\log\frac{1}{\epsilon}-\gamma_{0}\right)+W_{\Omega}(a,d;h_{\rm ex})+o(1)$ (2.14)
and
$\lim_{\epsilon\searrow 0}\lVert J_{A_{\epsilon}}(u_{\epsilon})-\pi\sum_{j=1}^{n}d_{j}\delta_{a_{j}}\rVert_{\dot{W}^{-1,q}}=0,$ (2.15)
for some $q\geq 1$, where $W_{\Omega}$ is the renormalized energy (2.9) and $\gamma_{0}$ is the constant defined in (2.12).
The main result of [KS11] is that such well-preparedness is preserved along time, which can be stated as follows:
Theorem 2.2 (Theorem 1.1 in [KS11]).
Let $(u_{\epsilon},A_{\epsilon},\Phi_{\epsilon}):\Omega\times[0,\infty)\rightarrow\mathbb{C}\times\mathbb{R}^{2}\times\mathbb{R}$ be a sequence of solutions of
$\begin{dcases}&\delta_{\epsilon}\left(\partial_{t}u_{\epsilon}+\imath\Phi_{\epsilon}u_{\epsilon}\right)=\nabla_{A_{\epsilon}}^{2}u_{\epsilon}+\frac{1}{\epsilon^{2}}u_{\epsilon}(1-|u_{\epsilon}|^{2}),\\ &\sigma_{\epsilon}\left(\partial_{t}A_{\epsilon}+\nabla\Phi_{\epsilon}\right)=-\operatorname{curl}h_{\epsilon}+(\imath\,u_{\epsilon},\nabla_{A_{\epsilon}}u_{\epsilon}),\end{dcases}\quad\mbox{on $\Omega\times\mathbb{R}_{+}$.}$ (2.16)
with $h_{\epsilon}=\operatorname{curl}A_{\epsilon}$ and boundary conditions
$\begin{dcases}&\nu\bm{\cdot}\nabla_{A_{\epsilon}}u_{\epsilon}=0,\\ &\nu\bm{\cdot}(\partial_{t}A_{\epsilon}+\nabla\Phi_{\epsilon})=0,\\ &h_{\epsilon}=h_{\rm ex},\end{dcases}\quad\mbox{on $\partial\Omega$}$ (2.17)
where the initial conditions $\{(u_{\epsilon}^{0},A_{\epsilon}^{0})\}_{\epsilon>0}$ are well-prepared with respect to some $a\in\Omega^{n}_{\ast}$ and $d\in\{\pm 1\}^{n}$. Moreover, we assume that
$\delta_{\epsilon}=\frac{\alpha_{0}}{|\log\epsilon|}+\imath\beta_{0}\quad\mbox{and}\quad\frac{\sigma_{0}}{|\log\epsilon|}\leq\sigma_{\epsilon}=o(1),$
where $\alpha_{0},\beta_{0}\geq 0$ and $\sigma_{0}>0$. Then for any time $t\leq T_{\max}$, where $T_{\max}>0$ denotes the maximal time of existence of the solution $a(t):[0,T_{\max})\rightarrow\Omega^{n}_{\ast}$ of the ODE
$\begin{dcases}\left(\alpha_{0}-d_{j}\beta_{0}\mathbb{J}\right)\dot{a}(t)=-\frac{1}{\pi}\nabla_{a_{j}}W_{\Omega}\left(a(t),d;h_{\rm ex}\right)\quad\mbox{for $t>0$},\\ a(0)=a,\end{dcases}\quad\mbox{where}\quad\mathbb{J}=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$, (2.18)
it holds that $\left(u(t),A(t)\right)$ are well-prepared with respect to $(a(t),d)$, i.e.,
  1. (i)
    (Jacobian) The gauged Jacobian $J_{A(t)}(u(t))$ satisfies
    $\lim_{\epsilon\searrow 0}\lVert J_{A_{\epsilon}(t)}(u_{\epsilon}(t))-\pi\sum_{j=1}^{n}d_{j}\delta_{a_{j}(t)}\rVert_{\dot{W}^{1,p}}=0,\quad\mbox{for any $1\leq p<2$.}$
  2. (ii)
    (Propagation of energy bound) We have
    $GL_{\epsilon}\left(u_{\epsilon}(t),A_{\epsilon}(t);h_{\rm ex}\right)=n\left(\pi\log\frac{1}{\epsilon}-\gamma_{0}\right)+W_{\Omega}(a(t),d;h_{\rm ex})+o(1).$
Remark (High applied magnetic fields).
While in [KS11], the authors consider $h_{\rm ex}=\mathcal{O}(1)$, in [SS04] the authors consider magnetic fields of order $h_{\rm ex}=\mathcal{O}(|\log\epsilon|)$ in the overdamped regime $\alpha=1$ and $\beta=0$. In this case, however, the magnetic field interaction dominates so that the limiting vortex dynamics has no vortex-vortex interaction.
Remark (The Schrödinger and heat flows).
Strictly speaking, only the case $\alpha_{0},\beta_{0}>0$ are treated in [KS11]. The case $\alpha_{0}>0$ and $\beta_{0}=0$ is addressed in [Spi02], and the case $\alpha_{0}=0$ and $\beta_{0}>0$ can be found in [Spi03].
As a corollary of Theorem 2.2, one can also show that the super-current and the magnetic field converge to their limiting counterparts. This result is not explicitly stated in [KS11, Theorem 1.1] but follows from their proof, see [KS11, Section 2.3].
Corollary 2.3 (Convergence of magnetic field).
Under the assumptions of Theorem 2.2, let $A_{\epsilon}$ be the solution satisfying the Coulomb gauge and $h_{\epsilon}=\mathrm{curl}\,A_{\epsilon}$, then we have
$\lim_{\epsilon\searrow 0}\lVert A_{\epsilon}(t)-A_{\ast}(t)\rVert_{\mathrm{L}^{p}(\Omega)}=0,\quad\mbox{and}\quad\lim_{\epsilon\searrow 0}\lVert h_{\epsilon}(t)-h_{\ast}(t)\rVert_{\mathrm{L}^{p}(\Omega)}=0,$
for any $0<t<T_{\max}$ and $1\leq p<\infty$ where $h_{\ast}\coloneqq-\Delta\Xi$, $A_{\ast}\coloneqq\mathrm{curl}\,\Xi$, and $\Xi$ is the solution of (2.11).
Remark 2.4 (Canonical harmonic map and convergence of super current).
Similarly, it follows from [KS11, Theorem 4.1] that the super-current of the sequence of solutions $u_{\epsilon}(t)$ converge to the super-current of the canonical harmonic map with vortex configuration $(a(t),d)$. More precisely, the canonical harmonic map with vortex configuration $(a,d)$ is the unique (up to a constant phase factor) $\mathbb{S}^{1}$-valued map $u_{\ast}$ satisfying
$\mathrm{curl}\,j(u_{\ast})=2\pi\sum_{j=1}^{n}d_{j}\delta_{a_{j}},\quad\mathrm{div}\,j(u_{\ast})=0,\quad\mbox{and}\quad\partial_{\nu}u_{\ast}=0\quad\mbox{on $\partial\Omega$}.$
Thus from [KS11, Theorem 4.1, eq. (4.8)], we have
$\lim_{\epsilon\searrow 0}\,\lVert j_{A_{\epsilon}}(u_{\epsilon})(t)-j_{A_{\ast}}(u_{\ast})(t)\rVert_{\mathrm{L}^{4/3}(\Omega)}=0,\quad\mbox{for any $0<t<T_{\max}$.}$
3. The limiting vortex dynamics
We now describe the procedure to numerically simulate the limiting ODE dynamics
$\begin{dcases}\left(\alpha_{0}-d_{j}\beta_{0}\mathbb{J}\right)\dot{a}(t)=-\frac{1}{\pi}\nabla_{a_{j}}W_{\Omega}\left(a(t),d;h_{\rm ex}\right)\quad\mbox{for $t>0$},\\ a(0)=a,\end{dcases}$ (3.1)
which is the central step in our method.
The starting point of this procedure is the following explicit formula for the gradient of $W_{\Omega}(a,d;h_{\rm ex})$. This formula is not immediate from (2.9) but a rigorous derivation can be found in [Spi03, Appendix A.2].
Lemma 3.1 (Gradient of renormalized energy).
Let $a\in\Omega^{n}_{\ast}$, $d\in\mathbb{Z}^{n}$ and $W_{\Omega}(a,d,h_{\rm ex})$ be the renormalized energy defined in (2.9). Then we have
$\nabla_{a_{k}}W_{\Omega}(a,d;h_{\rm ex})=-2\pi\sum_{j\neq k}d_{j}d_{k}\frac{a_{k}-a_{j}}{|a_{k}-a_{j}|^{2}}-2\pi d_{k}\nabla R(a_{k})-2\pi d_{k}\nabla\Xi(a_{k}),$ (3.2)
where $R:\Omega\rightarrow\mathbb{R}$ is the unique solution of
$\begin{dcases}\Delta R=0\quad&\mbox{in $\Omega$,}\\ R(x)=-\sum_{j=1}^{n}d_{j}\log|x-a_{j}|,\quad&\mbox{for $x\in\partial\Omega$,}\end{dcases}$ (3.3)
and $\Xi:\Omega\rightarrow\mathbb{R}$ is the unique solution of
$\quad\begin{dcases}\Delta^{2}\Xi-\Delta\Xi=2\pi\sum_{j=1}^{n}d_{j}\delta_{a_{j}},\quad&\mbox{in $\Omega$,}\\ \Xi=0,\quad&\mbox{on $\partial\Omega$,}\\ \Delta\Xi=-h_{\rm ex},\quad&\mbox{on $\partial\Omega$.}\end{dcases}$ (3.4)
Note that formula (3.2) requires, at each time step, the solution of two linear PDEs: a Laplace problem for the function $R$ defined in (2.10) and a modified biharmonic equation for the function $\Xi$ defined in (2.11). As thoroughly explained in [CKMS25], one can efficiently solve the Laplace problem by projecting the boundary conditions on the space of harmonic polynomials. However, as shown later in Lemma 3.4, it turns out that we do not need to compute $R$ to evaluate the gradient of $W_{\Omega}$. Therefore, let us focus now on the numerical scheme for solving the modified biharmonic equation (3.4).
3.1. The modified biharmonic equation
To solve the modified biharmonic equation (3.4), we first transform it into a homogeneous equation with nonhomogeneous boundary conditions. For this, let us recall (see, e.g., [HJ18, Equation 18]) that a fundamental solution of the operator $\Delta^{2}-\Delta$ is given by
$G(x,y)=G_{\Delta-1}(x,y)-G_{\Delta}(x,y)=-\frac{1}{2\pi}(K_{0}(r)+\log(r)),$ (3.5)
with r=|xy|r=|x-y| and where $K_{0}$ is the modified Bessel function of the second kind of order 0. This can be easily seen from the fact that $\Delta^{2}-\Delta=\Delta(\Delta-1)=(\Delta-1)\Delta$ and therefore
$-(\Delta^{2}-\Delta)\frac{1}{2\pi}(\log(r)+K_{0}(r))$ $=\Delta(\Delta-1)\left(\frac{-K_{0}(r)}{2\pi}\right)-(\Delta-1)\Delta\left(\frac{1}{2\pi}\log(r)\right)$
$=\Delta(\delta_{y})-(\Delta-1)(\delta_{y})=\delta_{y},$
where we used the fact that $\frac{1}{2\pi}\log(r)$ and $\frac{1}{2\pi}K_{0}(r)$ are, respectively, fundamental solutions of $\Delta$ and $\Delta-1$.
With the fundamental solution at hand, we can remove the singular source term in (3.4) at the cost of adding a smooth boundary term. More precisely, if we define
$\Xi_{p}(x)\coloneqq 2\pi\sum_{j=1}^{n}d_{j}\left(G\ast\delta_{a_{j}}\right)(x)=-\sum_{j=1}^{n}d_{j}(K_{0}+\log)(|x-a_{j}|)\quad\mbox{for $x\in\mathbb{R}^{2}$},$ (3.6)
then the solution of (3.4) is given by $\Xi=\Xi_{h}+\Xi_{p}$ where $\Xi_{h}$ solves
$\begin{dcases}\Delta^{2}\Xi_{h}-\Delta\Xi_{h}=0\quad&\mbox{in $\Omega$},\\ \Delta\Xi_{h}=-h_{\rm ex}-\Delta\Xi_{p},\quad&\mbox{on $\partial\Omega$,}\\ \Xi_{h}=-\Xi_{p}\quad&\mbox{on $\partial\Omega$.}\end{dcases}$ (3.7)
The function $\Xi_{h}$ can be further decomposed as follows.
Lemma 3.2 (Homogeneous solution).
Let $\Xi_{h}$ be the unique solution of (3.7) in $\mathrm{H}^{2}(\Omega)$, then we have
$\Xi_{h}=F-R+h_{\rm ex}$
where $R$ and $F$ are respectively the unique solutions in $\mathrm{H}^{1}(\Omega)$ of (3.3) and
$(\Delta-1)F=0,$ $\text{in }\Omega,$ (3.8a)
$F=-h_{\rm ex}+\sum_{j=1}^{n}d_{j}K_{0}(|x-a_{j}|),$ $\text{on }\partial\Omega.$ (3.8b)
Remark 3.3 (Uniqueness for the modified biharmonic equation in $\mathrm{H}^{2}(\Omega)$).
Note that the boundary trace of $\Delta\Xi_{h}$ is a priori not well-defined for merely $\mathrm{H}^{2}(\Omega)$ functions. However, (3.7) can be weakly formulated as follows: find $\Xi_{h}\in\mathrm{H}^{2}(\Omega)$ such that $\Xi_{h}=-\Xi_{p}$ in $\partial\Omega$ and
$\int_{\Omega}\Delta\Xi_{h}(\Delta-1)\psi=-\int_{\partial\Omega}\partial_{\nu}\psi(x)\left(h_{\rm ex}+\Delta\Xi_{p}(x)\right)\mathcal{H}^{1}(\mathrm{d}x),\quad\mbox{for any $\psi\in\mathrm{H}^{1}_{0}(\Omega)\cap\mathrm{H}^{2}(\Omega)$,}$ (3.9)
where $\partial_{\nu}\psi$ denotes the Neumann trace of $\psi$ along $\partial\Omega$. The existence and uniqueness of solutions of (3.9) then holds in any bounded Lipschitz domain $\Omega$.
Proof.
Let $F\in\mathrm{H}^{1}(\Omega)$ be a weak solution of (3.8) and $R\in\mathrm{H}^{1}(\Omega)$ be a weak solution of (3.3). Since the boundary functions are smooth, it follows from standard elliptic regularity that $F,R\in\mathrm{H}^{2}(\Omega)$. Thus, let us define $\widetilde{\Xi}_{h}=F-R+h_{\rm ex}\in\mathrm{H}^{2}(\Omega)$. Then in the distributional sense we have
$\Delta(\Delta-1)\widetilde{\Xi}_{h}=\Delta(\Delta-1)F-(\Delta-1)\Delta R=0\quad\mbox{in $\Omega$.}$
Since
$\Delta\widetilde{\Xi}_{h}=\Delta F=F,$ (3.10)
we have $\Delta\widetilde{\Xi}_{h}\in\mathrm{H}^{2}(\Omega)$ and therefore its boundary values are well-defined. Moreover, since
$\Delta\Xi_{p}=\sum_{j=1}^{n}d_{j}2\pi\Delta\left(G_{\Delta-1}-G_{\Delta}\right)\ast\delta_{a_{j}}=\sum_{j=1}^{n}d_{j}2\pi G_{\Delta-1}\ast\delta_{a_{j}}=-\sum_{j=1}^{n}d_{j}K_{0}(|x-a_{j}|),$ (3.11)
from the boundary condition for $F$, we see that
$\Delta\widetilde{\Xi}_{h}\rvert_{\partial\Omega}=F\rvert_{\partial\Omega}=-h_{\rm ex}+\sum_{j=1}^{n}d_{j}K_{0}(|x-a_{j}|)=-h_{\rm ex}-\Delta\Xi_{p}.$
Similarly, since
$\Xi_{p}(x)=-\sum_{j=1}^{n}d_{j}\left(\log(|x-a_{j}|)+K_{0}(|x-a_{j}|)\right),$ (3.12)
we have
$\widetilde{\Xi}_{h}\rvert_{\partial\Omega}=F\rvert_{\partial\Omega}-R\rvert_{\partial\Omega}+h_{\rm ex}=\sum_{j=1}^{n}d_{j}\left(K_{0}(|x-a_{j}|)+\log(|x-a_{j}|)\right)=-\Xi_{p}.$
Hence, $\widetilde{\Xi}_{h}$ solves (3.7) (and (3.9)) and therefore $\widetilde{\Xi}_{h}=\Xi_{h}$ by uniqueness of the solution. ∎
3.2. Simplified formulas for gradient, renormalized energy and magnetic field
We now use the decomposition in Lemma 3.2 to simplify some of the equations presented before. These simplified versions are the ones implemented in our algorithm.
We start with a simplified formula for the gradient of $W_{\Omega}$. This formula shows that only the function $F$ is necessary to simulate the ODE dynamics.
Lemma 3.4 (Simplified gradient).
Let $a\in\Omega^{n}_{\ast}$, $d\in\{\pm 1\}^{n}$, and $F$ be the solution of (3.8). Then
$\nabla_{a_{k}}W_{\Omega}(a,d;h_{\rm ex})=-2\pi\sum_{j\neq k}d_{j}d_{k}\frac{a_{k}-a_{j}}{|a_{k}-a_{j}|}K_{1}(|a_{j}-a_{k}|)-2\pi d_{k}\nabla F(a_{k}),$ (3.13)
where $K_{1}$ is the modified Bessel function of the second kind of order $1$.
Proof.
Recalling (3.12), we see that
$\nabla\Xi_{p}(x)=-\sum_{j=1}^{n}d_{j}\frac{x-a_{j}}{|x-a_{j}|}\left(\frac{1}{|x-a_{j}|}+\dot{K}_{0}(|x-a_{j}|)\right),\quad\mbox{for $x\not\in a$.}$
Next, note that by well-known identities for Bessel functions (see, e.g., [GR07, Section 8.446 and equation 18 in 8.486])
$\dot{K}_{0}(r)=-K_{1}(r),\quad\mbox{and}\quad\lim_{r\searrow 0}K_{1}(r)-\frac{1}{r}=0.$
Therefore,
$\nabla\Xi_{p}(a_{k})$ $=-\lim_{x\rightarrow a_{k}}\sum_{j=1}^{n}d_{j}\frac{x-a_{j}}{|x-a_{j}|}\left(\frac{1}{|x-a_{j}|}-K_{1}(|x-a_{j}|)\right)$
$=-\sum_{j\neq k}d_{j}\frac{a_{k}-a_{j}}{|a_{k}-a_{j}|}\left(\frac{1}{|a_{k}-a_{j}|}-K_{1}(|a_{k}-a_{j}|)\right).$ (3.14)
As the solution of (3.4) is given by $\Xi=\Xi_{p}+\Xi_{h}$, where $\Xi_{h}=F-R+h_{\rm ex}$ by Lemma 3.2, we conclude from (3.2) and (3.14) that
$\nabla_{a_{k}}W_{\Omega}(a,d;h_{\rm ex})$ $=-2\pi\sum_{j\neq k}d_{j}d_{k}\frac{a_{k}-a_{j}}{|a_{k}-a_{j}|^{2}}-2\pi d_{k}\nabla\Xi_{p}(a_{k})-2\pi d_{k}\nabla\Xi_{h}(a_{k})-2\pi d_{k}\nabla R(a_{k}),$
$=-2\pi\sum_{j\neq k}d_{j}d_{k}\frac{a_{k}-a_{j}}{|a_{k}-a_{j}|}K_{1}(|a_{k}-a_{j}|)-2\pi d_{k}\nabla F(a_{k}).$
Similarly, we can show that only the function $F$ is necessary to evaluate the renormalized energy.
Lemma 3.5 (Renormalized energy).
Let $(a,d)\in\Omega_{\ast}^{n}\times\mathbb{Z}^{d}$ and $F$ be the solution of (3.8), then the renormalized energy of $(a,d)$ can be written as
$W_{\Omega}(a,d;h_{\rm ex})=\pi\sum_{i\neq j}d_{i}d_{j}K_{0}(|a_{i}-a_{j}|)-\pi\sum_{j=1}^{n}d_{j}\left(F(a_{j})+h_{\rm ex}-d_{j}(\log(2)-\gamma)\right)+\frac{h_{\rm ex}^{2}}{2}|\Omega|\\ +\frac{h_{\rm ex}}{2}\int_{\Omega}\left(F(x)-\sum_{j=1}^{n}d_{j}K_{0}(|x-a_{j}|)\right)\mathrm{d}x,$ (3.15)
where $K_{0}$ is the modified Bessel function and $\gamma\approx 0.57721$ is the Euler-Mascheroni constant.
Proof.
The proof follows from a straigthforward calculation by plugging the following identities in (2.9): $\Xi+R=F+\Xi_{p}+h_{\rm ex}$,
$\int_{\partial\Omega}\partial_{\nu}\Xi=\int_{\Omega}\Delta\Xi=\int_{\Omega}F+\Delta\Xi_{p},$
equation (3.11), and
$\Xi_{p}(a_{j})=-\lim_{x\rightarrow a_{j}}\sum_{i=1}d_{i}(K_{0}+\log)(x)=-d_{i}\left(\log(2)-\gamma\right)-\sum_{i\neq j}d_{i}(K_{0}+\log)(|a_{i}-a_{j}|),$
where the last identity follows from the well-known asymptotics $K_{0}(r)=-\log(r)+\log(2)-\gamma+o(1)$ as $r\searrow 0$ (see [GR07, Section 8.443]). ∎
Lastly, we also obtain a simple formula for the magnetic field.
Lemma 3.6 (Magnetic field).
Let $\Xi$ be the solution of (3.4) for some $(a,d)\in\Omega_{\ast}^{n}\times\mathbb{Z}^{n}$, then the magnetic field $h_{\ast}=-\Delta\Xi$ is given by
$h_{\ast}(x)=-F(x)+\sum_{j=1}^{n}d_{j}K_{0}(|x-a_{j}|),$
where $F$ is the solution of (3.8).
Proof.
This is immediate from the identities $\Xi=\Xi_{p}+F-R+h_{\rm ex}$, $\Delta R=0$, $\Delta F=F$, and (3.11). ∎
3.3. Discretization of the modified biharmonic equation
As previously shown, to compute the solution of the modified biharmonic equation (3.4), it suffices to solve the harmonic equation (3.3) for $R$ and the modified Helmholtz equation (3.8) for $F$. To this end, we shall use the following functions:
$P_{j}(r,\theta)\coloneqq r^{|j|}\mathrm{exp}(\imath j\theta)\quad\mbox{and}\quad Q_{j}(r,\theta)\coloneqq I_{|j|}(r)\mathrm{exp}(\imath j\theta),\quad j\in\mathbb{Z},$ (3.16)
where $I_{|\alpha|}$ denotes the modified Bessel function of the first kind of order $|\alpha|$, and $(r,\theta)$ denote the polar coordinates centered at a suitable point $x_{0}\in\Omega$. More precisely, we use the harmonic polynomials $\{P_{j}\}_{j\in\mathbb{Z}}$ to approximate solutions of $R$ and the Bessel functions $\{Q_{j}\}_{j\in\mathbb{Z}}$ to approximate solutions for $F$.
Note that, from the defining equation for the Bessel function and the polar coordinate representation of the Laplacian, i.e.,
$r^{2}\partial_{r}^{2}I_{\alpha}+r\partial_{r}I_{\alpha}-(r^{2}+\alpha^{2})I_{\alpha}=0\quad\mbox{and}\quad\Delta=\partial_{r}^{2}+\frac{1}{r}\partial_{r}+\frac{1}{r^{2}}\partial_{\theta}^{2},$
it is easy to see that each $Q_{j}$ solves the modified Helmholtz equation $\Delta Q_{j}-Q_{j}=0$ in the whole $\mathbb{R}^{2}$. Similarly, $P_{j}$ is harmonic in $\mathbb{R}^{2}$. In particular, the restrictions $P_{j}\rvert_{\Omega}$ and $Q_{j}\rvert_{\Omega}$ to any open subset $\Omega$ are classical solutions of the respective equations in $\Omega$. We can therefore project the boundary conditions on the spaces spanned by those functions to compute approximations for $R$ and $F$.
For this, we shall work with the $\mathrm{L}^{2}(\Omega)$-orthogonal projections. More precisely, for a given discretization parameter $m\in\mathbb{N}$, we denote by $\mathbb{P}_{m}$ and $\mathbb{Q}_{m}$ respectively the $\mathrm{L}^{2}(\partial\Omega)$-orthogonal projection on the spaces
$H_{m}\coloneqq\mathrm{span}\{P_{j}\rvert_{\partial\Omega}:|j|\leq m\}\quad\mbox{and}\quad B_{m}\coloneqq\mathrm{span}\{Q_{j}\rvert_{\partial\Omega}:|j|\leq m\}$ (3.17)
For sufficiently regular domains (e.g. with piecewise smooth boundaries), such projections can be evaluated to numerical accuracy via suitable quadrature rules along the boundary $\partial\Omega$. Our numerical approximations $R_{m}$ and $F_{m}$ then correspond to the natural extension of the projected functions, i.e.,
$R_{m}(x)\coloneqq\sum_{|j|\leq m}a_{j}^{(m)}P_{j}(x)\quad\mbox{and}\quad F_{m}(x)\coloneqq\sum_{|j|\leq m}b_{j}^{(m)}Q_{j}(x),\quad\mbox{for $x\in\Omega$,}$ (3.18)
[…]
$\sum_{|j|\leq m}a_{j}^{(m)}P_{j}\rvert_{\partial\Omega}=\mathbb{P}_{m}\left(-\sum_{j=1}^{n}d_{j}\log|\cdot-a_{j}|\right)$ (3.19)
$\sum_{|j|\leq m}b_{j}^{(m)}Q_{j}\rvert_{\partial\Omega}=\mathbb{Q}_{m}\left(-h_{\rm ex}+\sum_{j=1}^{n}d_{j}K_{0}(|\cdot-a_{j}|)\right).$ (3.20)
It should be noted that the coefficients $\{a_{j}^{(m)},b_{j}^{(m)}\}_{|j|\leq m}$ may depend on $m$ as the boundary restrictions of $\{P_{j}\}_{|j|\leq m}$ or $\{Q_{j}\}_{|j|\leq m}$ are, in general, not $\mathrm{L}^{2}(\partial\Omega)$-orthogonal. Moreover, a priori, it is not clear that the functions $F_{m}$ and $R_{m}$ are able to approximate $F$ and $R$ with arbitrary accuracy or that the coefficients can be uniquely determined. However, the following result shows that this is indeed the case.
Theorem 3.7 (Basis property).
Let $\Omega\subset\mathbb{R}^{2}$ be a bounded simply connected domain with Lipschitz boundary. Then the families $\{P_{j}\rvert_{\partial\Omega}\}_{j\in\mathbb{Z}}$ respectively $\{Q_{j}\rvert_{\partial\Omega}\}_{j\in\mathbb{Z}}$ centered at any point $x_{0}\in\Omega$ are linearly independent and their spans are dense in $\mathrm{H}^{\frac{1}{2}}(\partial\Omega)$.
The key ingredient in the proof of Theorem 3.7 is the following Runge approximation property, whose proof can be found in [KV22, Section 3].
Lemma 3.8 (Theorem 3.9 in [KV22]).
Let $L=\Delta$ or $L=\Delta-1$, then for any Lipschitz open sets $\Omega\subset\subset\Omega^{\prime}$ such that $\Omega^{\prime}\setminus\overline{\Omega}$ is connected, any weak solution $u\in\mathrm{H}^{1}(\Omega)$ of $Lu=0$ in $\Omega$, and any $\epsilon>0$, there exists a weak solution $u^{\prime}\in\mathrm{H}^{1}(\Omega^{\prime})$ of $Lu^{\prime}=0$ in $\Omega^{\prime}$ such that
$\lVert u-u^{\prime}\rVert_{\mathrm{H}^{1}(\Omega)}<\epsilon.$
Proof of Theorem 3.7.
As the proof is essentially the same for both families of functions, let us focus only on $\{Q_{j}\}_{j\in\mathbb{Z}}$. We first show that the functions $\{Q_{j}\rvert_{\partial\Omega}\}_{j\in\mathbb{Z}}$ are linearly independent. For this, suppose that $Q_{j}=\sum_{k\neq j}\alpha_{k}Q_{k}$ in $\partial\Omega$ for finitely many non-zero $\alpha_{k}$. Then $Q\coloneqq Q_{j}-\sum_{k\neq j}\alpha_{k}Q_{k}$ solves $\Delta Q-Q=0$ and $Q\in\mathrm{H}^{1}_{0}(\Omega)$. As this operator is strictly negative, we must have $Q=0$ a.e. in $\Omega$. However, this is not possible for the following reason. As the $I_{\alpha}(r)$ are analytic functions for $r>0$, for any finite index set $\mathcal{I}\subset\mathbb{N}$, we can find a $\delta>0$ small enough such that $I_{k}(\delta)\neq 0$ for any $k\in\mathcal{I}$. Consequently, the functions $Q_{k}\rvert_{\partial B_{\delta}(x_{0})}=I_{|k|}(\delta)\mathrm{exp}(\imath k\theta)$ are linearly independent along $\partial B_{\delta}(x_{0})$, which is contained in $\Omega$ for $\delta$ small enough. Note that this argument shows that $I_{|k|}(r)\neq 0$ for any $r>0$.
We now prove the density result, namely, that any $g\in\mathrm{H}^{\frac{1}{2}}(\partial\Omega)$ can be well approximated by linear combinations of $\{Q_{j}\rvert_{\partial\Omega}\}_{j\in\mathbb{Z}}$. Let $u$ be a solution to $Lu=0$ in $\Omega$, with $u=g\in H^{1/2}(\partial\Omega)$ on $\partial\Omega$. Let $R>0$ be so large that $\Omega\subset\subset B_{R}(x_{0})$ and $B_{R}(x_{0})\setminus\overline{\Omega}$ is connected (which is possible because $\Omega$ is simply connected and bounded). Then, note that by Lemma 3.8, for any $\epsilon>0$ we can find $u^{\prime}\in\mathrm{H}^{1}(B_{R}(x_{0}))$ such that $\Delta u^{\prime}-u^{\prime}=0$ in $B_{R}(x_{0})$ and
$\lVert u^{\prime}-u\rVert_{\mathrm{H}^{1}(\Omega)}<\epsilon$ (3.21)
We can now approximate $u^{\prime}$ in $\mathrm{H}^{1}(B_{R}(x_{0}))$. More precisely, since $Q_{j}\rvert_{B_{R}(x_{0})}=I_{|j|}(R)\mathrm{exp}(\imath j\theta)$ are nothing but Fourier modes up to a nonzero factor (by the previous argument), from standard Fourier theory we can find finitely many $\alpha_{k}\in\mathbb{C}$ such that $v\coloneqq\sum_{k}\alpha_{k}Q_{k}$ satisfies
$\lVert v-u^{\prime}\rVert_{\mathrm{H}^{\frac{1}{2}}(\partial B_{R}(x_{0}))}<\epsilon.$
As $v-u^{\prime}$ is a $\mathrm{H}^{1}(\Omega)$ weak solution of $(\Delta-1)(v-u^{\prime})=0$ in $B_{R}(x_{0})$, it follows from standard elliptic regularity that
$\lVert v-u^{\prime}\rVert_{\mathrm{H}^{1}(B_{R}(x_{0}))}\lesssim_{R}\lVert v-u^{\prime}\rVert_{\mathrm{H}^{\frac{1}{2}}(\partial B_{R}(x_{0}))}<\epsilon.$ (3.22)
Hence, from the standard Sobolev trace theorem and estimates (3.21) and (3.22), we have
$\lVert v-u\rVert_{\mathrm{H}^{\frac{1}{2}}(\partial\Omega)}\lesssim_{\Omega}\lVert v-u\rVert_{\mathrm{H}^{1}(\Omega)}\lesssim\lVert u-u^{\prime}\rVert_{\mathrm{H}^{1}(\Omega)}+\lVert u^{\prime}-v\rVert_{\mathrm{H}^{1}(B_{R}(x_{0}))}\lesssim\epsilon,$ (3.23)
which completes the proof. ∎
As a corollary of Theorem 3.7, we see that the numerical approximations converge to the exact solutions in the limit $m\rightarrow\infty$. This provides a theoretical justification for our approach.
Corollary 3.9 (Convergence in the large $m$ limit).
For any bounded open set $\Omega$ with Lipschitz boundary, the numerical approximations $F_{m}$ and $R_{m}$ defined in (3.18) satisfy
$\lim_{m\rightarrow\infty}\lVert R-R_{m}\rVert_{\mathrm{H}^{\frac{1}{2}}(\Omega)}=\lim_{m\rightarrow\infty}\lVert F-F_{m}\rVert_{\mathrm{H}^{\frac{1}{2}}(\Omega)}=0,$ (3.24)
where $F$ and $R$ are the exact solutions of (3.8) and (3.3).
Proof.
As $\mathrm{H}^{\frac{1}{2}}(\partial\Omega)$ is densely embedded in $\mathrm{L}^{2}(\partial\Omega)$, the span of the sets $\{Q_{j}\rvert_{\partial\Omega}\}_{j\in\mathbb{Z}}$ or $\{P_{j}\rvert_{\partial\Omega}\}_{j\in\mathbb{Z}}$ are dense in $\mathrm{L}^{2}(\partial\Omega)$. Consequently, we have
$\lim_{m\rightarrow\infty}\lVert R-R_{m}\rVert_{\mathrm{L}^{2}(\partial\Omega)}=\lim_{m\rightarrow\infty}\lVert F-F_{m}\rVert_{\mathrm{L}^{2}(\partial\Omega)}=0.$
The convergence for $R$ in (3.24) now follows from the following deep estimate (see [Dah77, Dah79] or [JK95, Theorems 5.2–5.4 and Corollary 5.5.])
$\lVert u\rVert_{\mathrm{H}^{\frac{1}{2}}(\Omega)}\lesssim\lVert g\rVert_{\mathrm{L}^{2}(\partial\Omega)},$
for the unique harmonic function $u\in\mathrm{H}^{\frac{1}{2}}(\Omega)$ satisfying $u=g$ on $\partial\Omega$ (in the non-tangential sense). For the convergence of $F_{m}$ we need a similar estimate for solutions of $\Delta u-u=0$. This estimate can be established via a perturbative argument. Precisely, let $u^{\prime}\in\mathrm{H}^{\frac{1}{2}}(\Omega)$ be the unique harmonic function satisfying $u^{\prime}=g\in\mathrm{L}^{2}(\partial\Omega)$ in the non-tangential sense, then from the classical elliptic theory we can find a unique $w\in\mathrm{H}^{1}_{0}(\Omega)$ such that $\Delta w-w=u^{\prime}$ and $\lVert w\rVert_{\mathrm{H}^{1}(\Omega)}\lesssim\lVert u^{\prime}\rVert_{\mathrm{L}^{2}(\Omega)}$. It then follows that $u\coloneqq u^{\prime}+w$ solves
$\Delta u-u=0,\quad u=g\quad\mbox{on $\partial\Omega$ and satisfies}\quad\lVert u\rVert_{\mathrm{H}^{\frac{1}{2}}(\Omega)}\lesssim\lVert g\rVert_{\mathrm{L}^{2}(\partial\Omega)}.$
Remark 3.10 (Pointwise convergence).
Standard elliptic regularity theory (see, e.g. [GT01]) shows that estimates (3.24) implies interior pointwise $C^{k}$ convergence, i.e.,
$\lim_{m\rightarrow\infty}\lVert R-R_{m}\rVert_{C^{k}(V)}=\lim_{m\rightarrow\infty}\lVert F-F_{m}\rVert_{C^{k}(V)}=0.$
for any $V$ compactly contained in $\Omega$ and $k\in\mathbb{N}$. In particular, this shows that the gradient of the renormalized energy can be exactly evaluated in the large $m$-limit.
Remark 3.11 (Explicit estimate vs general domains).
In the case of a disk $\Omega=B_{r}(x_{0})$, the boundary restrictions of $P_{j}$ and $Q_{j}$ are nothing but the Fourier modes $\mathrm{exp}(\imath j\theta)$. In particular, in this case one can obtain explicit convergence and error estimates, as done in [CKMS25]. Nevertheless, we emphasize that the previous theorem justifies the method for a very general class of domains $\Omega$.
3.4. Numerical simulation of the limiting ODE
Now that we have introduced a novel discretization method to solve the modified biharmonic equation required to compute the gradient of the renormalized energy, we end this section by presenting a few trajectories for various initial conditions and external magnetic fields $h_{\rm ex}$.
The computational domain is the unit disk $\Omega=B_{1}(0)$, with different initial conditions $a^{0}$ and $d$ for each case. For the unit disk, the projection operators $\mathbb{P}_{m}$ and $\mathbb{Q}_{m}$ introduced in Section 3.3 take simpler forms since, in this case, they both coincide with truncated Fourier series, as mentioned in Remark 3.11. Solving the modified Helmholtz equation (3.8) can thus be done as follows:
  1. (1)
    Choose a maximum degree $m$, and fix it once and for all. Then, for any $g\in L^{2}(\partial\Omega)$,
[…]
  2. (2)
[…]
    $0,2\pi)\ni\theta\mapsto g_{a}(e^{\imath\theta})\coloneqq-h_{\rm ex}+\sum_{j=1}^{n}d_{j}K_{0}(|(\cos\theta,\sin\theta)-a_{j}|)$
    up to order $m$, for instance using a Fast Fourier Transform (FFT).
  3. (3)
    Compute the (approximate) solution $F_{m}$ to (3.8) by expanding $\mathbb{Q}_{m}g_{a}$ inside the disc:
[…]
    which is a solution of (3.8a) satisfying (3.8b) approximately.
We are now capable of simulating some trajectories. In the limiting ODE (3.1), we fix $\alpha_{0}=0$, $\beta_{0}=1$ so that we are in the case of the Schrödinger flow. This ODE is solved with a 4th-order Runge–Kutta method (RK4, see e.g. [HWN93, Chapter II]) with time step $\delta t=10^{-3}$ up to some time $T=1.0$. The gradient of the renormalized energy is computed with the simplified formulation (3.13), where $F$ solves (3.8a)-(3.8b) and is approximated with $m=64$ Fourier modes on the boundary $\partial\Omega$. We consider 4 different initial conditions, each of them being used twice with $h_{\rm ex}=3$ and $h_{\rm ex}=-2$. Note that all the trajectories were obtained in a few seconds on a personal laptop: a significant improvement when compared to the computational cost of the complete PDE (2.16) for small $\epsilon$.
Such trajectories can already be analyzed from a physics perspective: in the case of vortices with positive degree, adding a positive magnetic field $h_{\rm ex}>0$ brings an additional counterclockwise rotation speed to the vortices (as can be seen from Case 1 without magnetic field and Case 1 with $h_{\rm ex}=3$ where the vortices span a shorter distance within the same time), while adding a negative magnetic field $h_{\rm ex}<0$ brings an additional clockwise rotation speed (as can be seen from Case 2 without magnetic field and Case 2 with $h_{\rm ex}=-2$). In the next section, we present how to recover an approximate solution to the PDE from the trajectories obtained in the singular limit.
Table 3.1. Initial configurations for simulation of the Hamiltonian dynamics.
Case $n$ $a^{0}$ $d$
Case 11 22 $\left((-0.5,0.0),(0.5,0.0)\right)$ $(1,1)$
Case 22 33 $\left((-0.5,0.0),(0.5,0.0),(0.0,0.0)\right)$ (1,1,1)(1,1,-1)
Case 33 22 $\left((-0.75,0.0),(0.75,0.0)\right)$ (1,1)(-1,1)
Case 44 44 $\left((-0.6,-0.6),(-0.6,0.6),(0.6,0.6),(0.6,-0.6)\right)$ (1,1,1,1)(1,-1,1,-1)
Figure 1. Trajectories of vortices in the singular limit $\epsilon\searrow 0$. Initial conditions are represented by dots and the different degrees used are specified for each trajectory. From top to bottom: cases 1 to 4. From left to right: $h_{\rm ex}=0$, $h_{\rm ex}=3$, $h_{\rm ex}=-2$.
In Figure 2, we illustrate the numerical convergence of the trajectories, both in time and in $m$, on the numerical scheme we used to generate the limiting trajectories for $h_{\rm ex}=3$ and various initial conditions: we use $d=\{-1,+1\}$ and $a^{0}=\{(-\ell,0),(\ell,0)\}$ with different values of $\ell$ to mimic vortices that are either far apart, close to each other or close to the boundary $\partial\Omega$. As expected, the observed convergence rates are compatible with a 4th order in time and spectral accuracy in $m$ independently of the inter-vortex distance. However, when the vortices are too close to each other, or too close to the boundary, the prefactors in the convergence rate increase, as proven in [CKMS25] without magnetic field.
Figure 2. Numerical illustration of the convergence of the numerical method to obtain the limiting trajectories, for $h_{\rm ex}=3$. (Left) The slope shows the expected 4th order in time accuracy. (Right) The plots are compatible with spectral accuracy in the number of Fourier modes on the boundary.
4. Numerical scheme and results
The goal of this section is to introduce a numerical strategy to recover, at the PDE level, an approximation of the solution to the magnetic Ginzburg-Landau equation (2.16) from the, finite dimensional, vortex trajectories. First, we start with some complementary remarks on the numerical construction of well-prepared states.
4.1. Complements on well-prepared states
We focus here on the (numerical) construction of well-prepared states $(u_{\epsilon}^{0},A_{\epsilon}^{0})_{\epsilon>0}$, in the sense of Definition 2.1. Looking for a radially symmetric solution to the (magnetic free) stationary Ginzburg-Landau equation in a ball, one finds that its radial profile is given by the function $f_{\epsilon,r_{0}}$ that satisfies the ODE
$\frac{1}{r}\left(rf_{\epsilon,r_{0}}^{\prime}(r)\right)^{\prime}-\frac{1}{r^{2}}f_{\epsilon,r_{0}}(r)+\frac{1}{\epsilon^{2}}\left(1-|f_{\epsilon,r_{0}}(r)|^{2}\right)f_{\epsilon,r_{0}}(r)=0,$ (4.1)
together with the boundary conditions
$f_{\epsilon,r_{0}}(0)=0\quad\text{and}\quad f_{\epsilon,r_{0}}(r_{0})=1.$ (4.2)
For finite $r_{0}$, this equation can be solved numerically with very high precision (e.g. using a nonlinear finite element solver), even for $\epsilon$ as small as $10^{-3}$, see Figure 3. Finally, note that, by a scaling argument, $f_{\epsilon,r_{0}}$ only depends on the ratio $r_{0}/\epsilon$, so that we may write only $f_{\epsilon}$, see [KS11, Section 5] for more details.
Figure 3. Numerical approximation of $f_{\epsilon,r_{0}}$.
Starting from the localized approximation $(r,\theta)\mapsto f_{\epsilon}(r)e^{\imath\theta}$ of a single vortex, one can then build the following initial condition for (2.16): given vortex positions $a^{0}=(a_{j}^{0})_{j=1,\dots,n}\in\Omega_{*}^{n}$ with degrees $d=(d_{j})_{j=1,\dots,n}\in\{\pm 1\}^{n}$, we consider
$u_{\epsilon}^{*}(x;a^{0},d)=u^{*}(x;a^{0},d)\prod_{j=1}^{n}f_{\epsilon}(|x-a_{j}^{0}|)=\exp(\imath\varphi(x))\prod_{j=1}^{n}\left(\frac{x-a_{j}}{|x-a_{j}|}\right)^{d_{j}}f_{\epsilon}(|x-a_{j}^{0}|),\quad x\in\overline{\Omega}$ (4.3)
where $u^{*}$ is the canonical harmonic map with vortex locations $a^{0}$ (see Remark 2.4), and the phase $\varphi$ is computed to match Neumann boundary conditions. This numerical strategy can be seen as an operator from the manifold $\{u^{*}(a,d),\ a\in\Omega_{*}^{n},\ d\in\{\pm 1\}^{n}\}$ to $H^{1}(\Omega,\mathbb{C})$ which smoothes out the canonical harmonic map $u^{*}(a,d)$ at the singularities locations. It is identical to the smoothing strategy that we used in the magnetic-free case [CKMS25]. In the magnetic case, it is sufficient to consider in addition $A_{\epsilon}=A_{*}=\operatorname{curl}\Xi$ where $\Xi$ solves (2.11) to obtain well-prepared states $(u_{\epsilon}^{*},A_{*})_{\epsilon>0}$, see [KS11, Lemma 5.8]. Using this smoothing strategy, we are able to reconstruct an approximate solution to the infinite dimensional PDE as explained in the next section.
4.2. Description of the method
Instead of directly solving (2.16), we propose to take advantage of the trajectories of the vortices in the singular limit $\epsilon\searrow 0$ as well as the properties of the smoothing procedure used to obtain well-prepared initial conditions above. This yields the following algorithm:
  1. (1)
    Define initial vortex positions $a^{0}\in\Omega_{*}^{n}$ and degrees $d\in\{\pm 1\}^{n}$. Set up the initial phase $\varphi$ such that the initial condition $u_{\epsilon}^{0}=u_{\epsilon}^{*}(a^{0},d)$ satisfies the homogeneous Neumann boundary conditions. Compute and store the radial function $f_{\epsilon}$.
  2. (2)
    Evolve $a(t)$ according to the ODE (3.1) up to some maximum time $T$.
  3. (3)
    At time $t>0$, build back first an approximation of the wave function from the vortex positions $a(t)$ as
    $u_{\epsilon}^{*}(t)=u_{\epsilon}^{*}(a(t),d)=u^{*}(\cdot;a(t),d)\prod_{j=1}^{n}f_{\epsilon}(|\cdot-a_{j}(t)|),$ (4.4)
    where $u^{*}(x;a(t),d)$ is the canonical harmonic map (see Remark 2.4) defined by
    $u^{*}(x;a(t),d)=\exp(\imath\varphi(x))\prod_{j=1}^{n}\left(\frac{x-a_{j}}{|x-a_{j}|}\right)^{d_{j}},$
    with $\varphi$ the unique zero-mean harmonic function solving
    $\begin{cases}\Delta\varphi=0\quad\text{in }\Omega,\\ \displaystyle\partial_{\nu}\varphi(x)=-\sum_{j=1}^{n}d_{j}\partial_{\nu}\theta(x-a_{j})=\sum_{j=1}^{n}d_{j}\partial_{\tau}\ln|x-a_{j}|\quad\text{on }\partial\Omega.\end{cases}$ (4.5)
    Here, $\theta$ denotes the angle of a 2D vector (seen as the argument of the associated complex number), and $\nu$ (resp. $\tau$) the outward (resp. tangential) normal vector. Computing $\varphi$ with the appropriate boundary conditions is required so that the approximate solution $u_{\epsilon}^{*}(t)$ satisfies homogeneous Neumann boundary conditions. This Laplace’s equation can be solved using the same harmonic polynomial basis as the one used to solve (3.3). However, note that the harmonic function $\varphi$ is defined only up to a constant: this implies that the reconstructed wave function $u_{\epsilon}^{*}$ is an approximation of $u_{\epsilon}$ only up to a constant phase. Nevertheless, we note that a global phase factor is nothing but a trivial gauge transformation (see (2.6)), and therefore, all physical quantities of interest such as the super-current, the magnetic field or the vorticity are independent of this phase.
    An approximation of the magnetic potential AA can also be recovered by computing $\Xi$ as $\Xi_{h}+\Xi_{p}$ (see eqs. (3.6) and (3.7)) and then computing $A_{*}=\operatorname{curl}\Xi$. The magnetic field $h_{*}$ can be reconstructed with $h_{*}=\operatorname{curl}A_{*}$ or directly evaluated using Lemma 3.6.
Finally, the method we propose can be summarized by the diagram in Figure 4.
initial conditions
$(a^{0},d)\in\Omega_{*}^{n}\times\{\pm 1\}^{n}$
$\xrightarrow[\text{ODE}]{\text{evolve}}$ $(a(t),d)\in\Omega_{*}^{n}\times\{\pm 1\}^{n}$
$\downarrow$ smoothing smoothing $\downarrow$
well-prepared state
$u_{\epsilon}^{0}=u_{\epsilon}^{*}(a^{0},d)$, and $A_{*}$
$\xrightarrow[\text{TDGL}]{\text{evolve}}$ $u_{\epsilon}(t)\approx u_{\epsilon}^{*}(a(t),d)$ and $h_{\epsilon}(t)\approx h_{*}=\operatorname{curl}A_{*}$
(up to a constant phase)
Figure 4. Diagram summarizing the numerical simulation of the TDGL equation (2.16) via vortex tracking.
4.3. Vortex positions as initial conditions
We now present some numerical simulations of approximate solutions to (2.16) in the regime of small $\epsilon$. The reference vortex trajectories are those from Section 3.4. The solution of (4.1) is approximated numerically with high accuracy using 1D finite elements with small mesh size $\delta r=10^{-5}$ and $r_{0}=0.1$. This is computed once and for all in an offline stage, before running the simulation.
Finally, the solution to Laplace’s equation (4.5) in the smoothing process at time $t>0$ is computed using harmonic polynomials of degree $m=128$. Here, the initial vortex positions $a^{0}$ are provided as input data and the initial phase $\varphi$ is computed in order to match the Neumann boundary conditions (see (4.3)). We present the numerical results obtained by the application of our numerical method to cases 1, 2, 3 and 4 from Section 3.4, for $\epsilon=0.03$ and display, for each of them, the modulus and the phase of $u_{\epsilon}^{*}$ as well as the magnetic field at different times $t$. Note that the boundary conditions of the magnetic field $h_{*}$ do match the constant value of $h_{\rm ex}$ and that, independently of the sign of $h_{\rm ex}$, vortices with a positive (resp. negative) degree induce a positive (resp. negative) magnetic field.
[…]
Figure 8. Case $2h_{\rm ex}=$-2 : Squared modulus and phase of $u_{\epsilon}^{*}(t)$, with amplitude of $h_{*}$ for different times $t$.
Figure 9. Case $3h_{\rm ex}=$3 : Squared modulus and phase of $u_{\epsilon}^{*}(t)$, with amplitude of $h_{*}$ for different times $t$.
[…]
4.4. Numerical investigation of the small $\epsilon$ limit
We focus now on Case 1, in the case $h_{\rm ex}=3$: there are two vortices with identical winding number +1+1 and initial position $(\pm 0.5,0)$. From Figure 1, we know that we expect, for small values of $\epsilon$, the vortices to move around the circle of radius $0.5$ and centered at $(0,0)$.
4.4.1. Computational framework
In order to compare the numerical method presented in Section 4.2 to the exact solution of the TDGL equation (2.16), we compute a reference solution $u_{\epsilon}(t)$ with the finite element solver Gridap.jl [BV20] and meshing tool GMSH [GR09]. The magnetic TDGL equation (2.16) is solved for initial conditions given by the vortex configuration from Case 1 and the smoothing procedure described in Section 4.1. We use a forward Euler time discretization scheme and follow a numerical scheme adapted from [GX23]. The mesh size is of order $\delta x=5\cdot 10^{-2}$ with local a priori refined mesh on the vortex trajectory with mesh size $\delta x=3.5\cdot 10^{-3}$, see Figure 13. The time step is $\delta t=10^{-5}$. The finite element solver is then used to obtain various reference solutions $u_{\epsilon}$ for $\epsilon\in\{0.07,0.1,0.13\}$, up to $T=1$. We also compute a reference solution for $\epsilon=0.05$, with a finer refinement. Note that the simulation with such a mesh can take several days for small values of $\epsilon$: for instance, in our case, the simulations for $\epsilon=0.05$ took about ten days to run with a non-optimized code on a small sized cluster.
The reconstructed approximation $u_{\epsilon}^{*}(t)$ from Section 4.2 is obtained from the trajectories in the limit $\epsilon\searrow 0$, with $r_{0}=0.3$ and $m=128$ for the reconstruction step. The Hamiltonian dynamics is solved numerically with a RK4 method, with time step $\delta t=10^{-5}$ and $m=128$. Note that this time, the simulation only takes a few seconds on a personal laptop.
Figure 13. Mesh of the unit disk used to obtain reference solutions.
4.4.2. Numerical illustration of the singular limit $\epsilon\searrow 0$
In order to formally compare the reference solution $u_{\epsilon}(t)$ and the reconstructed approximation $u_{\epsilon}^{*}(t)$, we evaluate and compare different physical quantities, namely: the vortex trajectories, the super-current, the induced magnetic fields and the order parameter.
First, by tracking its isolated zeros, for instance using one of the algorithms introduced in [DLVN25, KSDH23], we are able to localize the vortices of the reference solution $u_{\epsilon}(t)$ and plot the associated trajectories in Figure 14, in the spirit of the study realized in [BT14]. Then, we plot in Figure 15 the following quantities:
$\frac{\lVert j_{A_{\epsilon}}(u_{\epsilon})(t)-j_{A_{*}}(u_{\epsilon}^{*})(t)\rVert_{L^{\frac{4}{3}}}}{\lVert j_{A_{\epsilon}}(u_{\epsilon})(t)\rVert_{L^{\frac{4}{3}}}},\quad\frac{\lVert h_{\epsilon}(t)-h_{*}(t)\rVert_{L^{2}}}{\lVert h_{\epsilon}(t)\rVert_{L^{2}}},\quad\text{and}\quad\frac{\lVert u_{\epsilon}(t)-u_{\epsilon}^{*}(t)\rVert_{L^{2}}}{\lVert u_{\epsilon}(t)\rVert_{L^{2}}}.$
Despite the increasing errors along time, the smaller $\epsilon$, the better the approximation of these quantities. This motivates further such a scheme to reduce the computational time required to solve the TDGL equation in the regime of small $\epsilon$. One should keep in mind though that this approximation remains valid as long as the vortices stay far from each other and from the boundary and is not capable for instance of describing the annihilation of vortices and the resulting nonlinear phenomena.
Figure 14. Vortex trajectories in the limit case for different values of $\epsilon>0$ and for $\epsilon=0$ (right).
Figure 15. Error on different quantities up to time $T$ for case 1. (Top left) Error on vortices positions in $\Omega$. (Top right) $L^{\frac{4}{3}}$ norm of the error on the super-currents. (Bottom left) $L^{2}$ norm of the error between $u_{\epsilon}(t)$ and the reconstructed wave function $u_{\epsilon}^{*}(t)$. (Bottom right) $L^{2}$ norm of the error between $h_{\epsilon}(t)$ and the reconstructed magnetic field $h_{*}(t)$ .
Finally, one can exploit the numerical results to extrapolate convergence rates with respect to $\epsilon$. In Figure 16, we perform a linear regression on the (absolute) error on the magnetic field and the super-current. For the magnetic field, the linear regression is performed at four different times while for the super-current, given the noisy results observed in Figure 15, the linear regression is performed after averaging the errors over time. For the magnetic field, the regressions clearly show that the error increase in time, but also that the power of $\epsilon$ decreases in time. As for the super current, the regression seems to indicate a slower convergence than for the magnetic field, with a power of epsilon of about $1/2$.
Figure 16. Linear regression with respect to $\epsilon$ of the error on the magnetic field and the super-current.
5. Conclusion
In this work, we first presented a numerical method to simulate the limiting ODE dynamics of point-like magnetic vortices arising from the TDGL equations with constant magnetic fields. From these trajectories, we proposed a numerical method to recover approximate solutions of the TDGL equations in the regime of small but finite $\epsilon$ under a constant external magnetic field. This method avoids the use of very fine meshes to properly capture the vortices core, thereby allowing for significantly faster and accurate simulations of the physical quantities of interest such as the super current, the magnetic field, and the density of Cooper pairs. Remarkably, the method becomes more accurate the smaller $\epsilon$ is. However, the accuracy is only guaranteed as long as the vortex pairs are sufficiently separated. In particular, this prevents accurate simulations of some physical phenomena such as radiation waves emitted by vortex-antivortex annihilation, which is left for future work.
Acknowledgements
TC acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Project number 442047500 through the Collaborative Research Center "Sparsity and Singular Structures" (SFB 1481). The authors are also grateful to Huadong Gao and Xie Wen for providing the Python code of [GX23] for preliminary tests.